5/6 times what equals 1/6
Answer by pmesler(52)

To start, we need to translate this sentence into a mathematical sentence. When I see "what" that automatically means x.
So another way to say this is
5/6 multiplied by x = 1/6 or
5/6x = 1/6.
To isolate x we need to multiply both sides by the reciprocal of 5/6 which is 6/5. We do this so the numerator and denominators divide out and all you're left with is x on the left-hand side.
so,
(6/5)5/6x = 1/6(6/5).
x = 6/30 = 1/5.
Therefore, 5/6 times 1/5 equals 1/6.
Let's write this out mathematically to see if it's correct.
5/6 * 1/5 = 1/6
5/30 = 1/6
Let's reduce the left-hand side by dividing 5/30 by 5/5.

We get 1/6 = 1/6. It checks out so it is the correct answer.